Carla Vista Diesel owns the Fredonia Barber Shop. He employs 7 barbers and pays each a base rate of $1.200 per month. One ofDetermine net income, assuming 1.800 haircuts are given in a month. Net income / (Loss) $
Answer #1

Salary of barber = $1,200 per month

Salary of six barbers = 1,200 x 6

= $7,200

Salary of barbers - manager = 1,200+555

= $1,755

Advertisement = $300 per month

Rent = $1,010 per month

Fixed utilities = $170 per month

Magazines = $30 per month

Total fixed cost per month = Salary of six barbers + Salary of barbers - manager + Advertising + Rent + Utilities + Magazines

= 7,200+1,755+300+1,010+170+30

= $10,465

Barber supplies = $0.5 per haircut

Barber commission = $5.90 per haircut

Utilities = $0.10 per haircut

Variable cost per unit = Barber supplies+ Barber commission + Utilities

= 0.5+5.90+0.10

= $6.50

Income Statement
Revenue (1,800 x 13) 23,400
Total variable cost (1,800 x 6.50) -11,700
Contribution Margin $11,700
Total fixed cost -10,465
Net Income $1,235

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks